Through what potential difference must an electron pass to have a wavelength of 500 angstrom.


Grade:

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Calculate E=hc/wavelength

Now E=eV

Calculate potential difference...

vikas askiitian expert
13 years ago

lamda  = h/(2mVe)1/2

 

h = 6.6*10-34 , m = 9.1*10-31 , e = 1.6*10-19

 

lamda = wavelength = 500angstromes = 500*10-10m

 

after substituting all these values we get

V = 5.86*10-4 or 0.586 millivolts